Do you think that the volume

Do you think that the volume control on an audio amplifier is calibrated in intensity or intensity level? In other words, if you turn your stereo half way down, have you cut the power in half or cut the decibel level in half? Explain.
Answer #1

Volume control is calibrated in intensity level.

intensity is given as

I = Io 10db/10                 Io = reference intensity ,   db = intensity level

Power is given as

P = I/A = Io 10db/10 /A

when we make the intensity level half, the power does not become half


answered by: jdpo
Answer #2

Sound intensity level (SIL) or acoustic intensity level is the level (a logarithmic quantity) of the intensity of a sound relative to a reference value. It is denoted \(L_{\mu}\), expressed in \(\mathrm{dB}\), and defined \(\mathrm{by}^{[2]}\)

\(L_{I}=\frac{1}{2} \ln \left(\frac{I}{I_{0}}\right) \mathrm{Np}=\log _{10}\left(\frac{I}{I_{0}}\right) \mathrm{B}=10 \log _{10}\left(\frac{I}{I_{0}}\right) \mathrm{dB}\)

the unit measured is decibels. so if we reduce the sterio half down the decibel elvel is cut to half but not the power.

answered by: monkey